1. Show that the function f(x)= | cos 2x| is a periodic function



2. Find the local extreme value of (1/x)^x , if it exists.

1
Expert's answer
2021-11-15T19:50:22-0500

1)

Consider the graph of f(x)=cos 2x below



From which the graph of f(x)=|cos 2x| is



Hence, as shown in the figure above the function f(x)=|cos 2x| is a periodic function with period "\frac{π}{2}"

2)

Given f(x)="(\frac{1}{x})^x" ,x>0

"ln(f(x))=ln((\frac{1}{x})^x)"

"ln(f(x))=xln(\frac{1}{x})"

"ln(f(x))=-xln(\frac{1}{x})"

Differentiate to get;

"\frac{1}{f(x)}\cdot f'(x)=-(x\cdot\frac{1}{x}+ln\ x)"

"\implies f'(x)=-(\frac{1}{x})^x(1+ln\ x)"

For critical points f'(x)=0

"\implies -(\frac{1}{x})^x(1+ln\ x)=0"

ln x=-1

"x=e^{-1}"

Hence local maximum exists and occurs at "x=e^{-1}"

Value "f(e^{-1})=(\frac{1}{e^{-1}})^{e^{-1}}"

="e^{e^{-1}}"